# SP-Chain: Boosting Intra-Shard and Cross-Shard Security and Performance in Blockchain Sharding

### v1

@@ 5 Protocol Design
This section presents the detailed design of SP-Chain. We first describe our two-phase concurrent voting intra-shard consensus in Sec. 5.1. Based on the proposed consensus protocol, we describe the leader rotation mechanism in Sec. 5.2. Next, we describe how cross-shard transactions are efficiently and securely processed in Sec. 5.3. Finally, finish this section by briefly describing the shard reconfiguration in Sec. 5.4.
5.1 Intra-Shard Consensus
@@
Our intra-shard consensus mechanism has two major design points: (1) Based on the feature that each shard contains a small number of nodes and has a good network synchronization rate, we design an efficient synchronous voting consensus protocol with two-phase communication; (2) Based on the two-phase voting, we propose the concurrent voting scheme that converts the serially executed steps in traditional consensus into parallel.
Two-Phase Voting. Our two-phase voting consensus protocol is divided into: phase 1, transaction package and block broadcast; phase 2, voting and block insertion. It is essentially a synchronous consensus mechanism, which achieves optimal resiliency of 1/2 in each shard and hence, allows total resiliency of 1/3 [37]. Details of the mechanism are described as follows.
Transaction Package and Block Broadcast. In each slot, the current leader packs the transactions into a block, and broadcasts the block and a simple block digest to its committee. The digest of a block mainly contains the current block number and slot number, the block hash, and the leader’s identity. The digest is used to do some pre-validation of the block. Since the size of the block digest is small, the member will receive the digest first. As there is a definite delay upper bound in the synchronous network model, we define the broadcast delay upper bound of the block digest as $\triangle_{d}$, which means every member in the shard will receive the block digest within $\triangle_{d}$ after the leader broadcasts the block and the digest. We also define the broadcast delay upper bound of the block and digest as $\triangle_{b}$, meaning that each member will receive the block and the digest within $\triangle_{b}$ after the leader broadcasts them.
Our consensus mechanism does not require clock synchronization between nodes. Suppose that a member in a shard receives the block digest at $\delta_{i}<\triangle_{d}$ moment after the broadcast. At that moment, it will start its clock and wait for the broadcasted block. Note that there is no requirement for all members to synchronize network clock information, as clock synchronization is impractical for a decentralized system like blockchain. We only require that each node’s time flow rate is consistent, which is a reasonable assumption [2].
Voting and Block Insertion. In normal case, after a member waits for $\triangle_{b}$, it will vote on the block. Members will verify whether the transactions are valid and broadcast voting information to the network. We define the delay upper bound of broadcasting voting information as $\triangle_{v}$. Since the clocks between nodes are not synchronized, each node during the voting process needs to wait for $\triangle_{d}+\triangle_{v}$ to ensure that every member of the shard receives the voting information of other members. After the waiting time, if more than $k/2$ ($k$ is the number of nodes in each shard) of the votes are received to approve the block, the member commit the block to its local ledger. Otherwise, it will treat the block as invalid and enter the next time slot.
@@
Remarks. There are usually 3 rounds of communication (e.g., pre-prepare, prepare, commit) in previous consensus protocols [23, 17, 13, 38] under partial-synchronous network. In our proposed protocol, there are 2 rounds of communication. The first round is block broadcast, which is similar to the pre-prepare phase in previous consensus. The second round is voting, which can be seen as the compaction of the prepare and commit phase. The two-phase voting consensus protocol exploits the features of the synchronous network in each shard, such as the guaranteed delay upper bound and the high connectivity between honest nodes. In this synchronous network context, the proposed protocol reduces the communication overhead without compromising security. More importantly, our design leverages the fact that each shard consists of a small number of nodes. This feature keeps $\triangle_{d}$, $\triangle_{b}$ and $\triangle_{v}$ at a small value, which shortens the interval of each slot and improves the protocol efficiency.
Concurrent Voting. To shorten the delay of each slot and further improve system performance, we propose a concurrent voting scheme based on two-phase voting, shown in Fig. 2. In each slot, the consensus protocol can be generally divided into four serially executed steps: transaction package, block broadcast, voting, and block insertion. With concurrent voting mechanism, these steps are broken down and reorganized. Specifically, in slot $t$, when the leader is packing transactions for the newly proposed block $b_{t}$, the members in the meantime are voting for the last proposed block $b_{t-1}$ (if the leader in slot $t-1$ did not propose a block, the value should be $b_{t-2}$). Additionally, when the leader broadcasts block $b_{t}$, members insert block $b_{t-1}$ at the same time.
@@
Design Challenges. The concurrent voting brings additional design challenges. Details are described as follows.
@@
Remarks. Our concurrent voting is different from previous pipelined consensus protocols [20, 36], where they focus on the parallelization of the communication. In concurrent voting, we decouple the consensus process in a more fine-grained way, where both communication and computation are parallelized. Also noting that the computation power is usually assumed to be sufficient. Therefore, the network is the bottleneck, rather than computation. Hence, in practice, the transaction package and block insertion (computation) take less time than voting and block broadcasting, respectively. Finally, since each node’s clock is not synchronized, a node may receive a vote (or digest) before its waiting time $\triangle_{d}+\triangle_{v}$ (or $\triangle_{d}$) is over. In this case, the node verifies the received message while waiting for the waiting time to end.
Figure 2: Overview of concurrent voting.
5.2 Leader Rotation
@@
Chain-Based Randomness Generation. At the beginning of each slot, a new random number is calculated to elect the leader of that slot. Specifically, when a node starts slot $t$, it extracts the signature information from the latest confirmed block (e.g., $b_{t-2}$, as the consensus is not reached for $b_{t-1}$ due to concurrent voting). Each node uses the signature information and the current slot number $t$ as a seed, input the seed to a publicly known pseudo-random number generation function (i.e., the random oracle mentioned in Sec. 3.2). This function will uniformly map the value of the seed to one of the nodes. The selected person is the leader in slot $t$ and is responsible for generating block $b_{t}$.
Choice of Signature Information. The most crucial point in the above process is the choice of signature information. When the selected signature information is not biased, it can be ensured that the leader election is unbiased. For this reason, we design that when $l_{t}$ (leader of slot $t$) generates a block, it will sign the current slot number $t$ and leave the signature information $SIG_{l_{t}}(t)$ in the block header, shown in Fig. 3. In this way, each member can use the signature from the latest confirmed block (e.g., $SIG_{l_{t-2}}(t-2)$) as the seed to calculate the new leader. For example:
$l_{t}=F(SIG_{l_{t-2}}(t-2),t),$ (1)
where $F(\cdot)$ is the random oracle that uniformly maps the input to one of the leader candidates.
5.3 Cross-Shard Transactions
@@
We leverage the Merkle Tree to verify the correctness of a transaction. A straightforward idea is to attach a proof (i.e., a complete Merkle path) to each cross-shard transaction so that the destination shard can verify the validity of the cross-shard transaction, but this approach introduces lots of overhead. To reduce the overhead, our main idea is that, for the transactions in a block, we arrange them and construct Merkle Tree according to different shards. Based on the constructed Merkle Tree, the transactions sent to the same shard are attached with one pruned Merkle proof together. One such proof can be used to verify this batch of transactions simultaneously.
Construction of Merkle Tree. The transactions packed into a block are no longer randomly arranged and constructed into a Merkle Tree as traditional approaches. In SP-Chain, transactions sent to the same shard will be sorted first according to their transaction hash values, as shown in Fig. 3. The sorted transactions sent to the same shard will then be constructed into a Merkle Subtree. Different Merkle Subtrees (representing transactions sent to different shards) will be merged into a complete Merkle Tree, including transactions sent to all shards. The Merkle root of the complete Merkle Tree will be written into the block header and verified by members during consensus.
@@
Receipt and Verification of Cross-Shard Transactions. After receiving the messages mentioned above, the destination shard reconstructs the corresponding Merkle Subtree root, and then reconstructs the Merkle Tree root based on other received Merkle Subtree roots. After the reconstruction, any node in the destination shard can judge whether the received transactions are modified by comparing the reconstructed Merkle Tree root with the Merkle Tree root in the signed block header. To prevent the leader from forging shard members and signatures, a shard member table (see Sec. 5.4 for details) is maintained by each node. The table contains the valid public keys of all nodes in all shards. In this way, after receiving the header, the destination shard can check whether there is an illegal member’s signature by comparing the member table.
Figure 3: Structure overview of a block.
5.4 Shard Reconfiguration
The main components of our shard reconfiguration are similar to RapidChain [37], which includes: 1) Offline PoW to prevent Sybil attacks; 2) Epoch randomness generation; 3) Committee reconfiguration; 4) node fast initialization after joining the committee. The main difference between our design and theirs lies in the epoch randomness generation. The epoch randomness is used to solve the offline PoW and reshuffle shard members. In SP-Chain, for efficiency and security, we use our chain-based randomness to generate epoch randomness. Specifically, since the block headers are broadcasted (Sec. 5.3), the reference committee [37, 1] can collect the seed information in all confirmed blocks in the last epoch. The reference committee then XOR the seeds to obtain a new epoch seed and use it to generate the epoch random number. Each new node can request the randomness of this epoch as a fresh PoW puzzle. Additionally, during shard reconfiguration, node change information (join or leave) will be broadcast. According to the node change information, each shard generates a state block containing the shard member table (of all shards).
6 Security and Performance Analysis
@@ 6.1 Epoch Security
Ensuring Negligible Failure Probability. We should carefully choose the shard size to bound the failure probability of the system to be negligible. As will be explained in Sec. 7.1, our choice of shard size can limit the failure probability to be less than $2^{-20}\approx 9\cdot 10^{-7}$ (time-to-failure of more than 4580 years for one-day epoch), which is a wide-adapted probability threshold [9, 37].

### v2

@@ V Protocol Design
This section presents the detailed design of SP-Chain. We first describe our two-phase concurrent voting intra-shard consensus in Section V-A. Based on the proposed consensus protocol, we describe the leader rotation mechanism in Section V-B. Next, we describe how cross-shard transactions are efficiently and securely processed in Section V-C. Finally, finish this section by briefly describing the shard reconfiguration in Section V-D.
Figure 2: Overview of two-phase voting intra-shard consensus.
V-A Intra-Shard Consensus
@@
Our intra-shard consensus mechanism has two major design points: 1) Based on the feature that each shard contains a small number of nodes and has a good network synchronization rate, we design an efficient, BFT-typed synchronous voting consensus protocol with two-phase communication; 2) Based on the two-phase voting, we propose the concurrent voting scheme that converts the serially handled computation and communication steps in traditional consensus into parallel.
V-A1 Two-Phase Voting
Each consensus round in a shard consists of two phases under a synchronous network model: phase 1 – block proposal and broadcast, followed by phase 2 – voting and block commitment. Figure 2 illustrates the workflow for a single round. By leveraging network synchrony (known message delay bounds) and a small committee size per shard, this design reduces the communication rounds from the usual three (pre-prepare, prepare, commit in PBFT-style protocols) to two, while still tolerating up to 50% faulty nodes in each shard (which corresponds to the standard <1/3 byzantine fault tolerance across the network).
Phase 1: Block Proposal and Broadcast. At the start of a slot, the elected leader assembles pending transactions into a new block and broadcasts the block along with a short digest of the block to all other shard members. The block digest includes essential metadata (e.g. block height, slot number, leader ID) and a hash of the block, allowing members to quickly perform a pre-check before the full block arrives. Because the digest is small, it propagates faster: $\Delta_{d}$ denotes the maximum network delay for the digest, whereas $\Delta_{b}$ is the maximum delay for the full block. Upon receiving the digest (at some time $\delta\leq\Delta_{d}$ from the start of broadcast), each member starts a local timer and waits for the full block. No global clock sync is required – each node uses the digest’s arrival as a reference point for timing. By time $\Delta_{b}$, every honest member should have received the complete block from the leader. (If the block fails to arrive in time or is invalid, the protocol will reject that leader’s proposal and move to the next slot.)
Phase 2: Voting and Block Commitment. Once a member’s local timer reaches $\Delta_{b}$ after digest reception, it proceeds to validate the received block and then broadcasts its vote on the block to the shard. A vote contains the validator’s signature, the current slot number, and an indication of whether the block is considered valid. We let $\Delta_{v}$ be the maximum network delay for vote messages. Because different validators may start their timers at slightly different moments (up to $\Delta_{d}$ apart due to network latency of the digest), each validator waits an additional $\Delta_{d}+\Delta_{v}$ after sending its own vote to ensure receipt of votes from all other members. By the end of this voting phase, every member will have collected votes from a majority of the committee (when the leader and network behaved correctly). If more than half of the shard’s nodes voted to approve the block, the block is committed to the shard’s blockchain. Each node then officially adds the new block to its ledger. If the block fails to gather a majority of yes-votes (or if the leader’s proposal was never received), the block is rejected and the shard simply moves on to the next slot without committing it.
@@ V-A1 Two-Phase Voting
Remarks. There are usually 3 rounds of communication (e.g., pre-prepare, prepare, commit) in previous consensus protocols [30, 21, 13, 50] under partial-synchronous network. In our proposed protocol, there are 2 rounds of communication. The first round is block broadcast, which is similar to the pre-prepare phase in previous consensus. The second round is voting, which can be seen as the compaction of the prepare and commit phase. The two-phase voting consensus protocol exploits the features of the synchronous network in each shard, such as the guaranteed delay upper bound and the high connectivity between honest nodes. In this synchronous network context, the proposed protocol reduces the communication overhead without compromising security. More importantly, our design leverages the fact that each shard consists of a small number of nodes. This feature keeps $\Delta_{d}$, $\Delta_{b}$ and $\Delta_{v}$ at a small value, which shortens the interval of each slot and improves the protocol efficiency.
Figure 3: Overview of pipelined concurrent voting.
V-A2 Concurrent Voting
To maximize throughput, SP-Chain pipelines the above two phases across consecutive slots, shown in Figure 3. This means nodes can work on different phases of different blocks in parallel. For example, while the shard members are busy voting on the block from slot $t-1$ (Phase 2 for block $b_{t-1}$), the new leader for slot $t$ can concurrently begin packaging transactions for the next block $b_{t}$ (the start of Phase 1 for that round). Likewise, when the leader of slot $t$ later broadcasts block $b_{t}$, the members are simultaneously committing block $b_{t-1}$. This two-slot overlap ensures that there is almost no idle time: computation (block assembly and insertion) is done in parallel with communication (broadcasting and voting) in a continuous pipeline.
@@ V-A2 Concurrent Voting
Additional Challenges and Solutions. The concurrent voting brings additional security and efficiency challenges. Details are described as follows.
@@
Remarks. Our concurrent voting is different from previous pipelined consensus protocols [25, 47], where they focus on the parallelization of the communication. In concurrent voting, we decouple the consensus process in a more fine-grained way, where both communication and computation are parallelized. Also noting that the computation power is usually assumed to be sufficient. Therefore, the network is the bottleneck, rather than computation. Hence, in practice, the transaction package and block commitment (computation) take less time than voting and block broadcasting, respectively. Finally, since each node’s clock is not synchronized, a node may receive a vote (or digest) before its waiting time $\Delta_{d}+\Delta_{v}$ (or $\Delta_{d}$) is over. In this case, the node verifies the received message while waiting for the waiting time to end.
Theoretical Basis for Parameter Choices ($\Delta_{d}$, $\Delta_{b}$, $\Delta_{v}$). Under the synchronous network assumption (i.e., every message is delivered within a known maximum delay $\Delta$), we choose the intra-shard communication timeouts $\Delta_{d}$, $\Delta_{b}$, and $\Delta_{v}$ to be small multiples of this bound for each consensus phase (see Section VII-A for exact values). The theoretical rationale is that allocating on the order of one full worst-case network delay per sub-phase provides enough slack to absorb worst-case delay variance while ensuring all honest nodes receive the required digests, blocks, and votes before progressing. By waiting up to a couple of $\Delta$ intervals in each phase (e.g., a slightly longer window for the larger block broadcast, and around one $\Delta$ for smaller digest and vote messages), the protocol tolerates maximum network latency differences so that even the slowest honest participant’s message arrives in time. At the same time, keeping these delays a bit longer than $\Delta$ preserves consensus responsiveness, as nodes move to the next step as soon as the expected worst-case propagation time elapses, avoiding any unnecessary idle waiting. This careful setting of $\Delta_{d}$, $\Delta_{b}$, and $\Delta_{v}$ thus upholds consensus safety under the synchronous model (no node decides prematurely) while minimizing latency overhead, maintaining both reliability and efficiency in reaching agreement.
V-B Leader Rotation
@@
Chain-Based Randomness Generation. At the beginning of each slot, a new random number is calculated to elect the leader of that slot. Specifically, when a node starts slot $t$, it extracts the signature information from the latest confirmed block (e.g., $b_{t-2}$, as the consensus is not reached for $b_{t-1}$ due to concurrent voting). Each node uses the signature information and the current slot number $t$ as a seed, input the seed to a publicly known pseudo-random number generation function (i.e., the random oracle mentioned in Section III-B). This function will uniformly map the value of the seed to one of the nodes. The selected person is the leader in slot $t$ and is responsible for generating block $b_{t}$.
Choice of Signature Information. The most crucial point in the above process is the choice of signature information. When the selected signature information is not biased, it can be ensured that the leader election is unbiased. For this reason, we design that when $l_{t}$ (leader of slot $t$) generates a block, it will sign the current slot number $t$ and leave the signature information $SIG_{l_{t}}(t)$ in the block header. In this way, each member can use the signature from the latest confirmed block (e.g., $SIG_{l_{t-2}}(t-2)$) as the seed to calculate the new leader. For example:
$l_{t}=H(SIG_{l_{t-2}}(t-2),t),$ (1)
where $H(\cdot)$ is the random oracle that uniformly maps the input to one of the leader candidates.
V-C Cross-Shard Transactions
@@
We leverage the Merkle Tree (i.e., a cryptographic tree structure where each leaf is a hash of data, and each non-leaf node is the hash of its children) to verify the correctness of a transaction. A straightforward idea is to attach a Merkle proof (i.e., a way to prove that a leaf is part of a Merkle tree by showing the necessary sibling hashes) to each cross-shard transaction so that the destination shard can verify the validity of the cross-shard transaction, but this approach introduces lots of overhead. To reduce the overhead, our main idea is that, for the transactions in a block, we arrange them and construct Merkle Tree according to different shards. Based on the constructed Merkle Tree, the transactions sent to the same shard are attached with one pruned Merkle proof together. One such proof can be used to verify this batch of transactions simultaneously. An illustration for processing cross-shard transactions is shown in Figure 4.
Constructing Merkle (Sub)trees According to Shards. The transactions packed into a block are no longer randomly arranged and constructed into a Merkle Tree as traditional approaches. In SP-Chain, transactions sent to the same shard will be sorted first according to their transaction hash values, as shown in Figure 4. The sorted transactions sent to the same shard will then be constructed into a Merkle Subtree. Different Merkle Subtrees (representing transactions sent to different shards) will be merged into a complete Merkle Tree, including transactions sent to all shards. The Merkle root of the complete Merkle Tree will be written into the block header and verified by members during consensus.
@@
Receipt and Verification of Cross-Shard Transactions. After receiving the messages mentioned above, the destination shard reconstructs the corresponding Merkle Subtree root, and then reconstructs the Merkle Tree root based on other received Merkle Subtree roots. After the reconstruction, any node in the destination shard can judge whether the received transactions are modified by comparing the reconstructed Merkle Tree root with the Merkle Tree root in the signed block header. To prevent the leader from forging shard members and signatures, a shard member table (see Section V-D for details) is maintained by each node. The table contains the valid public keys of all nodes in all shards. In this way, after receiving the header, the destination shard can check whether there is an illegal member’s signature by comparing the member table.
Figure 4: Cross-shard transactions processing in a block.
V-D Shard Reconfiguration
The main components of our shard reconfiguration are similar to RapidChain [48], which includes: 1) Offline PoW to prevent Sybil attacks; 2) Epoch randomness generation; 3) Committee reconfiguration; 4) node fast initialization after joining the committee. The main difference between our design and theirs lies in the epoch randomness generation. The epoch randomness is used to solve the offline PoW and reshuffle shard members. In SP-Chain, for efficiency and security, we use our chain-based randomness to generate epoch randomness. Specifically, since the block headers are broadcasted (Section V-C), the reference committee (a.k.a. beacon chain) [48, 42] can collect the seed information in all confirmed blocks in the last epoch. The reference committee then XOR the seeds to obtain a new epoch seed and use it to generate the epoch random number. Each new node can request the randomness of this epoch as a fresh PoW puzzle. Additionally, during shard reconfiguration, node change information (join or leave) will be broadcast. According to the node change information, each shard generates a state block containing the shard member table (of all shards).
VI Security and Performance Analysis
@@ VI-A Epoch Security
Ensuring Negligible Failure Probability. We should carefully choose the shard size to bound the failure probability of the system to be negligible. As will be explained in Section VII-A, our choice of shard size can limit the failure probability to be less than $2^{-20}\approx 9\cdot 10^{-7}$ (time-to-failure of more than 4580 years for one-day epoch), which is a wide-adapted probability threshold [8, 48].